OpenStudy (anonymous):

I need someone to explain this to me, I really don't understand... Using Baye's Theorem: It is known that 2% of the population has a certain allergy. A test correctly identifies people who have the allergy 98% of the time. The test correctly identifies people who do not have the allergy 94% of the time. A doctor decides that anyone who tests positive for the allergy should begin taking anti-allergy medication. Do you think this is a good decision? Why or why not? I know it's pretty long and a little bit of reading but it's on paper so I had to copy it, I wouldn't have if I didn't need the

OpenStudy (anonymous):

this is often very confusing one good way to approach this is to use some actual numbers and see what happens

OpenStudy (anonymous):

it makes "baye's theorem" far more undersantable

OpenStudy (anonymous):

Okay but how do I choose my numbers?

OpenStudy (anonymous):

i would pick a large number for the population, so that we get all whole number answers (although once we get the answer, we can redo it without using that crutch) lets say the population is 1000 people ok?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and if 2% of the population suffers from the allergies, how many people exactly will have it? i.e what is 2% of 1000 ?

OpenStudy (anonymous):

Umm...40? I'm not sure I did this right

OpenStudy (anonymous):

yes that is right

OpenStudy (anonymous):

oops sorry that is wrong

OpenStudy (anonymous):

dang it

OpenStudy (anonymous):

2% of 1,000 is \(.02\times 1000=20\)

OpenStudy (anonymous):

that is why i picked a nice round number like 1000 so the it would be easy to compute the percents

OpenStudy (anonymous):

That what I did but my calculator said i was 40...oh well what do I have to do next?

OpenStudy (anonymous):

ok so 20 people have the allergy, how many do not?

OpenStudy (anonymous):

By the way that already helps a lot

OpenStudy (anonymous):

20 out of 1000 which is... 20/10000=0.2%?

OpenStudy (anonymous):

hold on i think i have confused you

OpenStudy (anonymous):

we have a population of 1,000 2% have the allergy, so 20 have the allergy if out of 1,000 people, 20 have the allergy, how many (not what percent) do not have the allergy

OpenStudy (anonymous):

ooooh, sorry with the 10000 I got confused with another problem on my paper sorry. Umm so 980 don't have the allergy?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so we know how many have it, and how many do not now lets see how many people will test positive for it

OpenStudy (anonymous):

of the 20 people that have it, the test is 98% accurate what is 98% of 20?

OpenStudy (anonymous):

btw i hope it is clear that you do \(.98\times 20\)

OpenStudy (anonymous):

19.6?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

yeah I'm good with the percentage....most of the time

OpenStudy (anonymous):

i guess we have to work with the decimal, should have chosen a population of 10,000 then it would be whole numbers but too late now

OpenStudy (anonymous):

now we know \(980\) so NOT have the allergy, lets see how many of those will also test positive

OpenStudy (anonymous):

for them the test is 96% accurate so 96% of them will test negative, but that means 6% of them will test positive what is 6% of 980?

OpenStudy (anonymous):

oops bad math

OpenStudy (anonymous):

4% will test positive what is 4% of 980

OpenStudy (anonymous):

39.2?

OpenStudy (anonymous):

yeah i get that too

OpenStudy (anonymous):

ok so now how many people total (it is a decimal) will test positive?

OpenStudy (anonymous):

0.04 people? I'm confused

OpenStudy (anonymous):

we computed two numbers of people that test positive right? the ones with the allergy, 19.6 and the ones without the allergy, 39.2 what is the total?

OpenStudy (anonymous):

oh, 58.8?

OpenStudy (anonymous):

right, that is the total number that test positive

OpenStudy (anonymous):

out of those, we know that 19.6 actually have the allergy

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so the question is, IF you test positive, what is the probabilty you actually have the allergy that is the number of people who test positive and have the allergy, divided by the total number of people who test postive ie\[\frac{ 19.6}{58.8}\]

OpenStudy (anonymous):

or whatever you get when you write that as a decimal, i get \(0.337\) rounded, a little more than a third

OpenStudy (anonymous):

okay I get 0.33333etc.

OpenStudy (anonymous):

oh maybe i put the wrong numbers in yeah you are right

OpenStudy (anonymous):

one third in other words

OpenStudy (anonymous):

would you like to redo this using "baye's formula" ? will will do pretty much the same arithmetic, and get the same answer

OpenStudy (anonymous):

yeah, it would help a lot, my teacher just throws a formula and exercises at us

OpenStudy (anonymous):

ok do you have the formula you are supposed to use? i can make a guess if you like

OpenStudy (anonymous):

yeah I have it

OpenStudy (anonymous):

go ahead and write it, we can walk through it slowly

OpenStudy (anonymous):

It's \[P(A|B)=\frac{P(A|B)*P(A)}{ P(B)}\]

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

I have it right in front of me

OpenStudy (anonymous):

first off, it this is going to make sense, you need to know what \(P(A|B)\) means do you know what it means?

OpenStudy (anonymous):

oh crap your definition is wrong, but we will get to that in a second

OpenStudy (anonymous):

It means probability of A given B

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

and the definition (not baye's theorem) is \[P(A|B)=\frac{P(A\cap B)}{P(B)}\]

OpenStudy (anonymous):

My teacher made us study about a formula like that, but I really didn't get a thing

OpenStudy (anonymous):

baye's theorem, at least one form, is not quite what you wrote, the condition is switched on the right it is \[P(A|B)=\frac{P(B|A)*P(A)}{ P(B)}\]

OpenStudy (anonymous):

ok lets see if we can get it

OpenStudy (anonymous):

oh yeah my mistake

OpenStudy (anonymous):

ok now the numerator in your fraction \[P(B|A)P(A)\] is just another way to find \(P(A\cap B)\) the probability that both A and B occur

OpenStudy (anonymous):

That's where I get extremely confused

OpenStudy (anonymous):

lets use our example you want the probability you have the allergy GIVEN you test positive

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

we can put A = you have the allergy, B = you test positive then we are trying to find \(P(A|B)\) the probability you have the allergy given that you test positive

OpenStudy (anonymous):

let me know if i lost you, we need to go slow i am sure

OpenStudy (anonymous):

okay so far I think I'm okay...I think so for now

OpenStudy (anonymous):

ok to compute that we need the numerator \(P(A\cap B) \) i.e. the probability you have it AND test positive we were not told that number

OpenStudy (anonymous):

so far so good

OpenStudy (anonymous):

but we were told \(P(B)\) the probabily you have the disease, that was \(2\%=0.02\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and we were also told \(P(B|A)\) the probability you have the disease given that you test positive that was the \(98\%=0.98\)

OpenStudy (anonymous):

to find the numerator, multiply those together, i.e. the probability you have the disease AND test positive is the probability you have the disease times the probability you test positive GIVEN that you have the disease, i.e. \[.02\times.98=.0196\]

OpenStudy (anonymous):

you notice that is the same numerator we had before, just with the decimal moved over three places, because we did not start out with a population of 1,000 we just went straight for the number

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

Wait why did you change the 98% to a decimal. to make it easier to work with?

OpenStudy (anonymous):

when doing any work in math, you use numbers, not percents \(98\%=0.98\) as a number percents are nice to look at, but think about what you did when you wanted to find 98% of 20 you multiplied 20 by .98 right, not by 98

OpenStudy (anonymous):

ready to find the denominator?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

that is \(P(B)\) the probability you test positive remember how we found that before? we had to add two numbers, they were \(19.6\) and \(39.2\) they will be the same this time, but with the decimal in a different place

OpenStudy (anonymous):

the fist one is the \(.00196\) we already found, that is the number of people who test positive given that they have the disease

OpenStudy (anonymous):

so you divided it by 1000 right?

OpenStudy (anonymous):

we also need the number of people who test positive who do NOT have the disease

OpenStudy (anonymous):

well, yes, but we just found again computing \(.02\times .98=.00196\)

OpenStudy (anonymous):

to compute the probability that you test positive if you do not have the disease we used \(.04\times .980=.0392\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I'm sorry I really want to learn but I have to go to bed

OpenStudy (anonymous):

now add together to get the probability you test positive, that is \(.00196+.00392=.00588\)

OpenStudy (anonymous):

you did help a lot already though I'll make sure I have a lot of time next time I ask a question though

OpenStudy (anonymous):

yeah it is late you can look at this later and see that we are really using the same numbers as before

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

later

OpenStudy (anonymous):

later!